A patient with Asthma at one point has the following ABG’s: pH = 7.55, PaCO2= 30 mmHg, and PaO2= 65 mm Hg. What is accounting for these initial values? Later that patient, presents with the following ABG’s: pH = 7.35, PaCO2= 45 mmHg, and PaO2= 42 mm Hg. Some students have said that the patient is showing improvement, as pH and PaCO2 are closer to “normal” values. Why is this reasoning incorrect?
Answer #1

The First given values are pH = 7.55, Paco2= 30 mmHg and PaO2 = 65 mmHg. This values can be interpreted as Respiratory alkalosis.

Normal values are

pH= 7.35 - 7.45

paco2= 35- 45 mmHg

pao2 = 80 - 100 mmHg

The patient may have hyperventilation. Therefore excessive Co2 exhaled ( Decreased Co2 than normal) and Poor inhale of O2 (oxygen).

The second given values are pH= 7.35, PaCo2 = 45mm Hg, Pao2 = 42 mmHg.

Even though pH, Paco2 are within normal limit, there is extreme low o2 value. So there is no improvement in patient condition. The patient is in bad condition.